OpenStudy (anonymous):

(8+3)^2 + 4 * (-3) What is the value of this expression?

OpenStudy (ranga):

8 + 3 = 11 11^2 = 121 4*(-3) = -12 121 - 12 = ?

OpenStudy (anonymous):

8 + 3 = 11 11^2 = 121 4*(-3) = -12 121 - 12 = ? P.S. It's Me!